4. What is the value of


sum - infty ^ infty 1/{ (z+n)^ 2 +a^2}?


1
Expert's answer
2021-10-29T02:53:18-0400

"\displaystyle{\sum_{-\infin}^{\infin}}\frac{1}{(z+n)^2+a^2}=2\pi i\cdot \sum res\ f(z_0)" for "z_0" in the upper half-plane


 "z_0=\pm ia-n"

"res_1\ f(z_0)=\displaystyle{\lim_{z\to z_0}}[f(z)(z-z_0)]=\frac{1}{2ia}"


"res_2\ f(z_0)=-\frac{1}{2ia}" is not in the upper half-plane if a > 0


then:


"\displaystyle{\sum_{-\infin}^{\infin}}\frac{1}{(z+n)^2+a^2}=2\pi i/2ia=\pi/a"